OpenStudy (anonymous):

divide. 4/5 \ 2/3 A) 8/15 B)5/6 C)1 1/5 d)1 7/8

OpenStudy (australia.j.):

hold on

OpenStudy (retireed):

4/5 divided by 2/3 is equivalent to .... 4/5 times 3/2 multiply numerators and denominators and simplify

OpenStudy (australia.j.):

b

OpenStudy (retireed):

It's 4/5 divided by 2/3 wrong answer!

OpenStudy (australia.j.):

wasn't me. it was my friend she wants to help

OpenStudy (anonymous):

d?

OpenStudy (australia.j.):

it's a

OpenStudy (retireed):

Show me the work.

OpenStudy (australia.j.):

4 x 2 = 8 5 x 3 = 15

OpenStudy (australia.j.):

so it's 8/15

OpenStudy (australia.j.):

am i right

OpenStudy (retireed):

Your 0 for 2 try again.

OpenStudy (australia.j.):

how @retirEEd

OpenStudy (anonymous):

im positive its 1 1/5

OpenStudy (australia.j.):

wait a second you can't divide fractions.

OpenStudy (retireed):

4/5 divided by 2/3 is equivalent to .... 4/5 times 3/2 multiply numerators and denominators and simplify

OpenStudy (retireed):

Are you sure @peak?

OpenStudy (anonymous):

positive

OpenStudy (retireed):

What? wait a second you can't divide fractions. Another incorrect statement.

OpenStudy (anonymous):

i just did

OpenStudy (anonymous):

i just checked it on a calculater and its 1 1/5

OpenStudy (retireed):

I like the attitude @peak you are 100% correct. My delay was due to bad internet.

OpenStudy (anonymous):

k

OpenStudy (retireed):

Just for completeness... 4/5 times 3/2 = 12 / 10 = 1 and 2/10 which is 1 and 1/5 1 1/5